OpenStudy (ny,ny):

exponent question... mn^-4 _______ p^0q^-2

OpenStudy (phi):

you should remember that x^0 (or anything to the 0 power ) = 1

OpenStudy (ny,ny):

yeah.

OpenStudy (phi):

so you can simplify p^0 to 1 you can make the exponents positive if you "flip" the fraction in other words \[ \frac{1}{x^{-2}} = \frac{x^2}{1} \] and also \[ \frac{x^{-2}}{1} = \frac{1}{x^{2}}\]

OpenStudy (phi):

What is the exact question ?

OpenStudy (ny,ny):

\[mn ^{-4}\div p ^{0} q ^{-2}\]